Hi can you please help me solve for this problem from this book: Doane−Seward: Applied Statistics in Business and Economics,chapter 10:
A new drug is being tested to see if it reduces the number of migraine headaches. Assuming equal variances, at α = .025, do these samples of number of monthly migraines from eight volunteers who are taking the drug and eight who are not show a significant reduction in the mean number of monthly migraine headaches? (See Science News 165, no. 9 [February 28, 2004].)
Topiramate: 3 3 2 4 3 4 3 5

Control: 3 3 5 5 7 5 5 4
Thank you

Answer by stanbon(48568)

A new drug is being tested to see if it reduces the number of migraine headaches. Assuming equal variances, at α = .025, do these samples of number of monthly migraines from eight volunteers who are taking the drug and eight who are not show a significant reduction in the mean number of monthly migraine headaches? (See Science News 165, no. 9 [February 28, 2004].)
-------------
Topiramate: 3 3 2 4 3 4 3 5
u1 bar = 3.375
s1 = 0.916
-------------
Control: 3 3 5 5 7 5 5 4
u2 bar = 4.625
s2 = 1.302
----------------
Ho: u1-u2 = 0
Ha: u1 - u2 < 0
Unknown variances; assumed equal:
test statistic
pooled estimate of common variance: sp^2 = [6*0.016^2 + 6*1.302^2]/[14]= 0.05
t statistic: t = (3.375-4.625)/sqrt[0.05(1/8 + 1/8)] = -11.18
p-value: P(-20< t<-11.18 with df = 8+8-2 = 14) = 0.00000001416
---
Conclusion: Since p-value is less than alpha, reject Ho.
The medicine is more effective than the control.
===================================================
Cheers,
Stan H.